OpenStudy (anonymous):

3e ^x+1 - 2=0 solve for x

OpenStudy (anonymous):

start with \[3e^{x+1}=2\] then \[e^{x+1}=\frac{2}{3}\] write in equivalent logarithmic for to get \[x+1=\ln(\frac{2}{3})\] then subtract 1 to get \[x=\ln(\frac{2}{3})-1\]

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

yw